When Grace is as old as her father is now, her brother will be twice as old as he is now and the age Grace will be when her brother is as old as her father is now. The total age of the three is a century. How old is each person?
Thank you.

Answer by rwm(914)

g+x=f
b+x=2b
g+b+f=100